# Shedding Light on Gravity: Black Hole Shadows and Lensing Signatures in Lorentz Gauge Theory

### v1

@@ II Black holes in LGT
Overall, these derivations underscore how curvature invariants—constructed from the Weyl tensor and its derivatives—serve as robust tools for detecting horizons and exploring the geometrical properties of black hole spacetimes. They provide an alternative method to the traditional approach of locating horizons by solving $g^{rr}=0$, and offer additional insights into the interplay between the spacetime geometry and the parameters that define the gravitational field.
@@ III Calculation of Hawking Radiation in Jacobi Metric formalism
where S is the classical action. The key idea is that the tunneling probability is dominated by the exponential of the imaginary part of S.
The corresponding Jacobi metric, which encapsulates the effective geometry experienced by the particle, is given by Gibbons (2016); Chanda et al. (2017); Das et al. (2017):
$\mathrm{d}s^{2}=j_{ij},\mathrm{d}x^{i}\mathrm{d}x^{j}$
$=\Big{(}E^{2}-m^{2}f(r)\Big{)}\left(\frac{\mathrm{d}r^{2}}{f^{2}(r)}+\frac{r^{2}}{f(r)}\Big{(}\mathrm{d}\theta^{2}+\sin^{2}\theta,\mathrm{d}\phi^{2}\Big{)}\right).$ (15)
Here, E and m are the energy and mass of the tunneling particle, respectively, and $f(r)$ is the lapse function that encodes the gravitational potential. This metric effectively redefines the “distance” in configuration space in such a way that the kinetic term of the particle’s motion is modulated by the combination $E^{2}-m^{2}f(r)$.
In the context of Hawking radiation, the tunneling process occurs predominantly along the radial direction near the horizon. Thus, the action for the particle undergoing radial tunneling can be written as
$S=-\int\sqrt{j_{ij},\frac{\mathrm{d}x^{i}}{\mathrm{d}s},\frac{\mathrm{d}x^{j}}{\mathrm{d}s}},\mathrm{d}s~{}.$ (16)
@@
$\sqrt{j_{ij},\frac{\mathrm{d}x^{i}}{\mathrm{d}s},\frac{\mathrm{d}x^{j}}{\mathrm{d}s}}=\pm\left(E^{2}-m^{2}f(r)\right)^{\frac{1}{2}}f^{-1}(r)\frac{\mathrm{d}r}{\mathrm{d}s}~{}.$ (17)
@@
$p_{r}=\partial_{r}S=\mp\left(E^{2}-m^{2}f(r)\right)^{\frac{1}{2}}f^{-1}(r)~{}.$ (18)
Since the tunneling process is confined to the near-horizon region, the condition $f(r)<0$ is imposed. In our expression for $p_{r}$, the sign ensures that the outgoing particle (with $p_{r}>0$) corresponds to the appropriate tunneling direction, in line with conventional treatments such as those found in Srinivasan and Padmanabhan (1999).
Near the horizon, the full four-dimensional metric simplifies to an effective (1+1)-dimensional form because the angular components become subdominant Robinson and Wilczek (2005); Iso et al. (2006); Majhi (2010). To capture the physics near the horizon, we expand the lapse function $f(r)$ in a Taylor series about the horizon $r=r_{H}$:
@@
$\approx 2\tilde{\kappa}(r-r_{H})+\mathcal{O}((r-r_{H})^{2})~{},$ (19)
@@
$\tilde{\kappa}=\frac{1}{2}f’(r_{H})~{}.$ (20)
@@
$S=\mp\frac{E}{2\tilde{\kappa}}\int_{r_{H}-\epsilon}^{r_{H}+\epsilon}\frac{\mathrm{d}r}{(r-r_{H})}\pm\frac{m^{2}}{2E}\int_{r_{H}-\epsilon}^{r_{H}+\epsilon}\mathrm{d}r$
$\mp\int_{r_{H}-\epsilon}^{r_{H}+\epsilon}\mathcal{O}(r-r_{H}),\mathrm{d}r~{}$ (21)
@@
$\int_{r_{H}-\epsilon}^{r_{H}+\epsilon}\frac{\mathrm{d}r}{(r-r_{H})}=-i\pi~{}.$ (22)
@@
[…]
Here, the positive sign corresponds to the outgoing trajectory, while the negative sign corresponds to the incoming trajectory.
The WKB wave functions for the outgoing and incoming particles can now be written as
$\psi_{\rm out}=A,e^{\frac{i}{\hbar}S_{\rm out}},\quad\psi_{\rm in}=A,e^{\frac{i}{\hbar}S_{\rm in}},$ (24)
@@
$P_{\rm in}=|\psi_{\rm in}|^{2}=|A|^{2}e^{\frac{\pi E}{\hbar\tilde{\kappa}}}~{}.$ (26)
@@
$\Gamma=\frac{P_{\rm out}}{P_{\rm in}}=e^{-\frac{2\pi E}{\hbar\tilde{\kappa}}}\equiv e^{-\frac{E}{T_{H}}}~{}.$ (27)
@@
$T_{H}=\frac{\tilde{\kappa}}{2\pi}~{}.$ (28)
@@
$\tilde{\kappa}=\left.\sqrt{-\frac{1}{2}\nabla_{\mu}\chi_{\nu}\nabla^{\mu}\chi^{\nu}},\right|{r=r_{H}}=\left.\frac{1}{2}\frac{\partial f(r)}{\partial r},\right|{r=r_{H}}~{}.$ (29)
@@
$T_{H}=\frac{m_{p}^{2}}{A_{0}^{4}m_{s}}~{},$ (30)
which encapsulates the dependence on the parameters $m_{p}$, $m_{s}$, and the scale parameter $A_{0}$.
Figure 1 illustrates how the Hawking temperature $T_{H}$ varies with the black hole mass parameter $m_{s}$ for different values of $A_{0}$. Figure 1 shows that as $m_{s}$ increases, the temperature rapidly decreases, indicating that heavier black holes radiate less intensely. Additionally, smaller values of $A_{0}$ result in higher temperatures, implying that configurations with lower $A_{0}$ are thermodynamically less stable.
[…]
In summary, this semi-classical tunneling method based on the Jacobi metric provides a consistent framework to derive the Hawking temperature. The detailed derivation—from the form of the Jacobi metric through the evaluation of the tunneling action and the resulting probability—shows that the Hawking temperature is intimately linked to the surface gravity of the black hole. This method not only reinforces the interpretation of Hawking radiation as a quantum tunneling process but also highlights the interplay between the black hole’s geometric parameters and its thermodynamic properties.
@@ IV Weak Deflection Angle Using the Gauss-Bonnet Theorem
The Gaussian optical curvature associated with this metric, which measures the intrinsic curvature affecting photon trajectories, is calculated using the Ricci scalar R as a relation $\mathcal{K}=R/2$. Explicitly, Gaussian curvature becomes:
@@
where $\kappa$ represents the geodesic curvature of the boundary, $\theta_{i}$ is the exterior angle at the $i^{th}$ vertex, and $\chi(\mathcal{D}_{R})=1$ denotes the Euler characteristic for a simply connected domain.
We choose our region outside the photon trajectory, and since the photon path $\gamma_{\tilde{g}}$ is geodesic, its geodesic curvature vanishes, i.e., $\kappa(\gamma_{\tilde{g}})=0$. For the boundary at infinity, represented by the curve $C_{R}:=r(\varphi)=R=\text{constant}$, the geodesic curvature is computed via:
@@
thus combining Eqs. (36) and (37), we have $\kappa(C_{R})\mathrm{d}t=\mathrm{d}\varphi$.
Using the straight-line approximation for photon trajectories, the radial coordinate is related to the impact parameter $b$ as $r=b/\sin\varphi$. Consequently, the Gauss–Bonnet theorem simplifies considerably, providing a straightforward integral representation for the deflection angle Gibbons and Werner (2008):
@@
$\mathrm{d}S=\left(\frac{3A_{0}^{5}m_{s}}{8\pi m_{p}^{2}}+A_{0}^{3}r\right)\mathrm{d}r\,\mathrm{d}\varphi.$ (39)
@@
Figure 2: Weak deflection angle plotted as a function of radial distance parameter $r$ for various values of the scale parameter $A_{0}$. The Planck mass squared is fixed at $m_{p}^{2}=1$.
Figure 2 illustrates how the weak deflection angle depends on the impact parameter $b$ for different values of the parameter $A_{0}$. Notably, the deflection angle decreases gradually as one moves away from the black hole, underscoring how gravitational lensing diminishes at larger distances. Furthermore, varying $A_{0}$ significantly influences lensing strength, with smaller values yielding stronger gravitational deflection.
@@ IV.1 Derivation of deflection angle using geodesics method
which follow from symmetries of the spacetime and corresponding constants of motion $E$ and $L$.
Substituting these expressions into the null geodesic condition ($\mathcal{L}=0$), we have
$-f(r)\left(\frac{E}{f(r)}\right)^{2}+\frac{\dot{r}^{2}}{f(r)}+r^{2}\left(\frac{L}{r^{2}}\right)^{2}=0\,,$ (42)
@@
The above equation is the radial geodesic equation, expressing the radial velocity in terms of the energy and angular momentum constants of motion.
Introducing the inverse radial coordinate $u\equiv 1/r$, we note
@@
This equation explicitly relates the inverse radius $u$ and the azimuthal angle $\phi$.
Differentiating Eq. (49) again gives the second-order orbital equation:
@@
which governs the photon trajectory around the massive source.
To solve perturbatively, let $u(\phi)=u_{0}(\phi)+\delta u(\phi)$, where $u_{0}$ solves the unperturbed equation (with $m_{S}=0$) and $\delta u$ is a small correction.

### v2

@@ II Black holes in LGT
Overall, these derivations underscore how curvature invariants—constructed from the Weyl tensor and its derivatives—serve as robust tools for detecting horizons and exploring the geometrical properties of black hole spacetimes. They provide an alternative method to the traditional approach of locating horizons by solving $g^{rr}=0$, and offer additional insights into the interplay between the spacetime geometry and the parameters that define the gravitational field. The solution is not gauge-equivalent to the standard Schwarzschild solution. It represents a physically distinct spacetime. Two metrics are gauge-equivalent (i.e., describe the same physical spacetime) if one can be transformed into the other by a coordinate transformation. A definitive way to test this is to compare their scalar curvature invariants, such as the Kretschmann scalar ($K=R^{\alpha\beta\gamma\delta}R_{\alpha\beta\gamma\delta}$), which are independent of the coordinate system. If the invariants are different, the spacetimes are physically distinct. By inspection, $K_{\text{proposed}}\neq K_{\text{Schw}}$ unless $A_{0}=1$. If $A_{0}\neq 1$, the second term in $K_{\text{proposed}}$ is non-zero and has a different radial dependence, making it impossible to match $K_{\text{Schw}}$ for all $r$. The main points are as follows: One may try
$\tilde{t}=\frac{t}{A_{0}},\quad\tilde{r}=\frac{r}{A_{0}},\quad\tilde{m}_{S}=\frac{m_{S}}{A_{0}}\,$, (18)
under which the trt\!-\!r sector becomes $d\tilde{t}^{2}+\bigl{(}1-\tfrac{\tilde{m}_{S}}{4\pi m_{P}^{2}\tilde{r}}\bigr{)}^{-1}d\tilde{r}^{2}$. However, the angular part transforms as
$r^{2}\,d\Omega^{2}=A_{0}^{2}\,\tilde{r}^{2}\,d\Omega^{2}\;\neq\;\tilde{r}^{2}\,d\Omega^{2}\quad(\text{unless }A_{0}^{2}=1),$ (19)
showing that in $(t,r,\theta,\varphi)$ coordinates one cannot eliminate $A_{0}$ without distorting the two‐sphere. A direct computation gives for metric (1):
$R\;=\;\frac{2\,(1-A_{0}^{2})}{r^{2}},\quad R_{\rm Schw}=0.$ (20)
They coincide only if $A_{0}^{2}=1$, confirming that metric (1) is not the usual Schwarzschild form in these coordinates. Since a physical invariant differs, no coordinate transformation can map one metric to the other. Therefore, the solution is not merely a rescaled version of Schwarzschild; it describes a different gravitational field with measurably different curvature.
@@ III Calculation of Hawking Radiation in Jacobi Metric formalism
where S is the classical action. The key idea is that the tunneling probability is dominated by the exponential of the imaginary part of S. The corresponding Jacobi metric, which encapsulates the effective geometry experienced by the particle, is given by Gibbons (2016); Chanda et al. (2017); Das et al. (2017):
$\mathrm{d}s^{2}=j_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j}=\Big{(}E^{2}-m^{2}f(r)\Big{)}\left(\frac{\mathrm{d}r^{2}}{f^{2}(r)}+\frac{r^{2}}{f(r)}\Big{(}\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\phi^{2}\Big{)}\right).$ (22)
Here, $M$ and $m$ are the energy and mass of the tunneling particle, respectively, and $f(r)$ is the lapse function that encodes the gravitational potential. This metric effectively redefines the “distance” in configuration space in such a way that the kinetic term of the particle’s motion is modulated by the combination $E^{2}-m^{2}f(r)$. In the context of Hawking radiation, the tunneling process occurs predominantly along the radial direction near the horizon. Thus, the action for the particle undergoing radial tunneling can be written as
$S=-\int\sqrt{j_{ij}\frac{\mathrm{d}x^{i}}{\mathrm{d}s}\frac{\mathrm{d}x^{j}}{\mathrm{d}s}}\mathrm{d}s~.$ (23)
@@
$\sqrt{j_{ij}\frac{\mathrm{d}x^{i}}{\mathrm{d}s}\frac{\mathrm{d}x^{j}}{\mathrm{d}s}}=\pm\left(E^{2}-m^{2}f(r)\right)^{\frac{1}{2}}f^{-1}(r)\frac{\mathrm{d}r}{\mathrm{d}s}.$ (24)
@@
$p_{r}=\partial_{r}S=\mp\left(E^{2}-m^{2}f(r)\right)^{\frac{1}{2}}f^{-1}(r).$ (25)
Since the tunneling process is confined to the near-horizon region, the condition $f(r)<0$ is imposed. In our expression for $p_{r}$, the sign ensures that the outgoing particle (with $p_{r}>0$) corresponds to the appropriate tunneling direction, in line with conventional treatments such as those found in Srinivasan and Padmanabhan (1999). Near the horizon, the full four-dimensional metric simplifies to an effective (1+1)-dimensional form because the angular components become subdominant Robinson and Wilczek (2005); Iso et al. (2006); Majhi (2010). To capture the physics near the horizon, we expand the lapse function $f(r)$ in a Taylor series about the horizon $r=r_{H}$:
@@
$\approx 2\tilde{\kappa}(r-r_{H})+\mathcal{O}((r-r_{H})^{2}),$ (26)
@@
$\tilde{\kappa}=\frac{1}{2}f’(r_{H}).$ (27)
@@
$S=\mp\frac{E}{2\tilde{\kappa}}\int_{r_{H}-\epsilon}^{r_{H}+\epsilon}\frac{\mathrm{d}r}{(r-r_{H})}\pm\frac{m^{2}}{2E}\int_{r_{H}-\epsilon}^{r_{H}+\epsilon}\mathrm{d}r\mp\int_{r_{H}-\epsilon}^{r_{H}+\epsilon}\mathcal{O}(r-r_{H})\mathrm{d}r,$ (28)
@@
$\int_{r_{H}-\epsilon}^{r_{H}+\epsilon}\frac{\mathrm{d}r}{(r-r_{H})}=-i\pi.$ (29)
@@
$S=\pm\frac{i\pi E}{2\tilde{\kappa}}+\textrm{(real part)}.$ (30)
Here, the positive sign corresponds to the outgoing trajectory, while the negative sign corresponds to the incoming trajectory. The WKB wave functions for the outgoing and incoming particles can now be written as
$\psi_{\rm out}=Ae^{\frac{i}{\hbar}S_{\rm out}},\quad\psi_{\rm in}=Ae^{\frac{i}{\hbar}S_{\rm in}},$ (31)
@@
$P_{\rm in}=|\psi_{\rm in}|^{2}=|A|^{2}e^{\frac{\pi E}{\hbar\tilde{\kappa}}}$. (33)
@@
$\Gamma=\frac{P_{\rm out}}{P_{\rm in}}=e^{-\frac{2\pi E}{\hbar\tilde{\kappa}}}\equiv e^{-\frac{E}{T_{H}}}.$ (34)
@@
$T_{H}=\frac{\tilde{\kappa}}{2\pi}.$ (35)
@@
$\tilde{\kappa}=\left.\sqrt{-\frac{1}{2}\nabla_{\mu}\chi_{\nu}\nabla^{\mu}\chi^{\nu}}\right|_{r=r_{H}}=\left.\frac{1}{2}\frac{\partial f(r)}{\partial r}\right|_{r=r_{H}}.$ (36)
@@
$T_{H}=\frac{m_{p}^{2}}{A_{0}^{4}m_{s}},$ (37)
which encapsulates the dependence on the parameters $m_{p}$, $m_{s}$, and the scale parameter $A_{0}$. By considering the Schwarzschild mass, $M=m_{s}/(8\pi m_{p}^{2})$, and substitute $m_{s}=8\pi Mm_{p}^{2}$ into Eq. (37)
$T_{H}\;=\;\frac{1}{8\,\pi\,M\,A_{0}^{4}}\,.$ (38)
Figure 1 illustrates how the Hawking temperature $T_{H}$ varies with the black hole mass parameter $M$ for different values of $A_{0}$. Figure 1 shows that as $M$ increases, the temperature rapidly decreases. Additionally, smaller values of $A_{0}$ result in higher temperatures.
Figure 1: Hawking temperature plotted as a function of the mass parameter $m_{s}$ for various values of $A_{0}=0.5,\,1,\,2,\,3$. Here, we have set $m_{p}^{2}=1$.
In summary, this semi-classical tunneling method based on the Jacobi metric provides a consistent framework to derive the Hawking temperature. The detailed derivation—from the form of the Jacobi metric through the evaluation of the tunneling action and the resulting probability—shows that the Hawking temperature is intimately linked to the surface gravity of the black hole.
@@ IV Weak Deflection Angle Using the Gauss-Bonnet Theorem
The Gaussian optical curvature associated with this metric, which measures the intrinsic curvature affecting photon trajectories, is calculated using the Ricci scalar R as a relation $\mathcal{K}=R/2$. A straightforward calculation yields
$\mathcal{K}=\frac{1}{2}\left(f(r)f^{\prime\prime}(r)-\frac{1}{2}f^{\prime}(r)^{2}\right)$ (40)
where primes denote $d/dr$. Explicitly, Gaussian curvature becomes:
@@
where $\kappa$ represents the geodesic curvature of the boundary, $\theta_{i}$ is the exterior angle at the $i^{th}$ vertex, and $\chi(\mathcal{D}_{R})=1$ denotes the Euler characteristic for a simply connected domain. We choose our region outside the photon trajectory, and since the photon path $\gamma_{\tilde{g}}$ is geodesic, its geodesic curvature vanishes, i.e., $\kappa(\gamma_{\tilde{g}})=0$. For the boundary at infinity, represented by the curve $C_{R}:=r(\varphi)=R=\text{constant}$, the geodesic curvature is computed via:
@@
thus combining Eqs. (45) and (46), we have $\kappa(C_{R})\mathrm{d}t=\mathrm{d}\varphi$. Using the straight-line approximation for photon trajectories, the radial coordinate is related to the impact parameter $b$ as $r=b/\sin\varphi$. Consequently, the Gauss–Bonnet theorem simplifies considerably, providing a straightforward integral representation for the deflection angle Gibbons and Werner (2008):
@@
$\mathrm{d}S\approx\left(\frac{3A_{0}^{5}m_{s}}{8\pi m_{p}^{2}}+A_{0}^{3}r\right)\mathrm{d}r\,\mathrm{d}\varphi.$ (48)
@@
Using the definition of the Schwarzschild mass, $M=m_{s}/(8\pi m_{p}^{2})$, we substitute $m_{s}=8\pi Mm_{p}^{2}$ into Eq. (49). This yields the deflection angle in the desired form:
$\alpha\approx\frac{4A_{0}M}{b}+\frac{3\pi A_{0}^{3}M^{2}}{4b^{2}}$ (50)
Figure 2: Weak deflection angle plotted as a function of radial distance parameter $r$ for various values of the scale parameter $A_{0}$. The Planck mass squared is fixed at $M=1$.
Figure 2 illustrates how the weak deflection angle depends on the impact parameter $b$ for different values of the parameter $A_{0}$. We constrain the parameter $A_{0}$ by equating the theoretical prediction in Eq. (50) with high-precision observational data from light deflection at the solar limb. The observed angle, $\theta$, is related to the General Relativistic prediction, $\delta_{\rm GR}=4M_{\odot}/R_{\odot}\approx 1.7516687^{\prime\prime}$, by the PPN parameter $\gamma$:
$\theta=\frac{1+\gamma}{2}\delta_{\rm GR}$ (51)
The accepted value for $\gamma$ is $0.99992\pm 0.00012$ Lambert and Le Poncin-Lafitte (2011). Setting $\alpha=\theta$ for a light ray grazing the Sun ($M=M_{\odot}$, $b=R_{\odot}$), we have:
$A_{0}\delta_{\rm GR}+\frac{3\pi A_{0}^{3}}{16}\delta_{\rm GR}^{2}=\frac{1+\gamma}{2}\delta_{\rm GR}$ (52)
Given that $\delta_{\rm GR}\approx 8.492\times 10^{-6}$ rad, the term proportional to $\delta_{\rm GR}^{2}$ is negligible. The equation simplifies to $A_{0}\approx(1+\gamma)/2$. Using the central value of $\gamma$, we find $A_{0}\approx 0.99996$. The uncertainty is dominated by that of $\gamma$, giving $\Delta A_{0}\approx\Delta\gamma/2=0.00006$.
By re-expressing the deflection angle in terms of the Schwarzschild mass and comparing it with observational data, we derive a stringent constraint on the model parameter $A_{0}$:
$A_{0}=0.99996\pm 0.00006$ (53)
This result shows that $A_{0}$ is consistent with unity, as predicted by General Relativity, to a high degree of precision.
@@ IV.1 Derivation of deflection angle using geodesics method
which follow from symmetries of the spacetime and corresponding constants of motion $E$ and $L$. Substituting these expressions into the null geodesic condition ($\mathcal{L}=0$), we have
$f(r)\left(\frac{E}{f(r)}\right)^{2}+\frac{\dot{r}^{2}}{f(r)}+r^{2}\left(\frac{L}{r^{2}}\right)^{2}=0\,$, (55)
@@
The above equation is the radial geodesic equation, expressing the radial velocity in terms of the energy and angular momentum constants of motion. Introducing the inverse radial coordinate $u\equiv 1/r$, we note
@@
This equation explicitly relates the inverse radius $u$ and the azimuthal angle $\phi$. Differentiating Eq. (62) again gives the second-order orbital equation:
@@
which governs the photon trajectory around the massive source. To solve perturbatively, let $u(\phi)=u_{0}(\phi)+\delta u(\phi)$, where $u_{0}$ solves the unperturbed equation (with $m_{S}=0$) and $\delta u$ is a small correction.